# Shifting AI Efficiency From Model-Centric to Data-Centric Compression

### v2

@@ 4.2 Suboptimal Token Representation
Most existing token compression methods fall into two categories: redundancy-based approaches that maximize information preservation between original ($\mathbf{X}$) and compressed tokens ($\mathbf{X^{\prime}}$) via $\max_{\mathcal{C}}I(\mathbf{X};\mathbf{X^{\prime}})$, and importance-based methods that ensure predictive sufficiency through $I(\mathbf{X^{\prime}};\mathbf{Y})\geq I(\mathbf{X};\mathbf{Y})-\epsilon$ and $\epsilon$ is the bearable information loss. While effective in their respective objectives, we argue that these paradigms share a critical limitation: neither guarantees that the compressed tokens $\mathbf{X^{\prime}}$ form an optimal representation for downstream modeling. The redundancy-based framework, despite preserving maximal mutual information with $\mathbf{X}$, often retains tokens with reconstructive but low discriminative value. The importance-based framework, on the other hand, prioritizes maintaining predictive performance with respect to the target variable $\mathbf{Y}$, but often at the cost of introducing task-specific biases. By focusing solely on the preservation of information relevant to a predefined label, these methods may overlook the need to maintain stable structural and semantic patterns across the token sequence in $\mathbf{X}$ that could enhance generalization across diverse downstream tasks. Consequently, both approaches risk producing token representations that are misaligned with the ultimate goal of effective and generalizable downstream modeling.
@@ 4.3 Fair Comparison
Rethinking FLOPs and Compression Ratios as Efficiency Metrics. Many token compression methods report speedup by estimating FLOPs reductions or directly using token compression ratios. But do FLOPs or compression ratios truly reflect real acceleration? Our analysis shows that, even with similar compression ratios or FLOPs, methods often vary significantly in runtime latency. Investigating further, we find: (i) Importance-based compression often uses attention scores Chen2024:FastV ; zhang2024sparsevlm , but this can limit compatibility with efficient attention mechanisms like Flash Attention dao2022flashattention ; dao2023flashattention2 , potentially contributing to the substantial discrepancy between FLOPs and actual runtime latency observed in some approaches. (ii) Some methods pursue high compression via progressive compression across layers xing2024pyramiddrop , adding overhead that offsets the gains from token reduction. Thus, we argue that runtime latency should be prioritized in evaluations, as FLOP or token count reductions don’t always yield real-world speedup.
Figure 2: Empirical comparison of carefully designed token compression methods and random token dropping. Results demonstrate that in multiple scenarios, some meticulously engineered token compression methods surprisingly underperform compared to random token pruning.
Token Compression Evaluation: The Benchmarking Gap. Current token compression methods are primarily evaluated using general-purpose benchmarks, which are not tailored to capture the unique challenges of token compression. As a result, some benchmarks—such as ScienceQA lu2022learn and VizWiz bigham2010vizwiz —paradoxically show improved performance under certain compression settings, or minimal degradation across varying compression ratios. These observations defy intuition and suggest that existing benchmarks may fail to meaningfully reflect the trade-offs introduced by token compression. This discrepancy raises concerns about the validity of current evaluation practices. In particular, benchmarks that fail to penalize information loss introduced by compression may obscure meaningful differences between methods. Additionally, the lack of task diversity and compression-sensitive metrics further limits our understanding of how these methods behave in realistic scenarios. Without dedicated benchmarks, it remains unclear whether observed gains are due to genuine improvements in compression quality or artifacts of misaligned evaluation settings.
@@ 5.1 Data-Model Centric Compression Co-Development
5.2 Dedicated Benchmarks for Token Compression
@@
Given the current limitations in evaluating data-centric token compression methods using general-purpose benchmarks, we envision the development of a dedicated benchmark specifically designed to evaluate them. Such a benchmark should comprehensively span diverse domains—including natural language processing, computer vision, and multi-modal tasks—and incorporate task-specific challenges particularly relevant to token compression, such as optical character recognition (OCR) parsing ouyang2024omnidocbench ; zhang2024ocr and automatic speech recognition (ASR) conneau2023fleurs ; park2024long . Furthermore, it is essential that this benchmark jointly considers both task performance and latency, as both are critical for real-world deployment. A well-rounded benchmark of this nature would enable more rigorous, fair, and holistic evaluation of token compression techniques, ultimately driving progress in this area.
6 Alternative Positions
@@
6.1 Model-Centric Compression as a Superior Alternative
@@
Model-centric compression methods, such as pruning Network_Pruning ; JMLR:v12:huang11b ; Deep_Compression ; Rao2021:DynamicViT ; jang2017Gumbel-Softmax ; Pan2021:IA-RED2 , quantization rokh2023comprehensive ; Wang:HAQ ; zhou2018adaptive ; yang2019quantization ; Lin:AWQ ; Frantar:GPTQ , and knowledge distillation Hinton:KD ; zhang2019self_distillation ; zhang2021self_distillation ; park2019relational_distillation , have long been established as effective techniques for reducing model size and computational cost. Proponents argue that this paradigm is reliable for deployment in resource-constrained environments and maintains performance consistency. For example, pruning techniques such as DynamicViT Rao2021:DynamicViT dynamically remove uninformative tokens during inference, reducing the computational load by up to 30–40% with minimal impact on accuracy. Proponents of this view claim that this approach achieves substantial speedups without discarding any original data. In contrast, data-centric methods that prune input tokens risk removing critical contextual information, which may degrade performance.
@@
Counterargument. Although model-centric compression is effective, it faces scalability issues as models and datasets grow, requiring costly full retraining and processing of entire inputs. In contrast, data-centric compression reduces input complexity upfront, easing computational burdens. Some data-centric methods update only a small parameter subset Zhao2024:DyT ; liu2024Sparse-Tuning ; Lee2024:DETM , while others enable training-free deployment bolya2023tome ; Chen2024:FastV ; li2024snapkv . Combining both approaches can improve efficiency without sacrificing accuracy azeemi2023data ; kousar2025pruning , making data-centric methods a complement to model-centric techniques.
6.2 Advanced Model Architectures as a more Promising Direction
@@
Another argument against data-centric compression is the continued advancement of model architectures that can inherently handle large datasets and long sequences more efficiently gu2021efficiently ; goldstein2024goldfinch ; peng2023rwkv . The development of transformer-based architectures, such as Vision Transformers ViT , Swin Transformers liu2021swin , and large language models like GPT-3 brown2020language , has shown significant improvements in both accuracy and scalability. These architectures integrate advanced techniques, such as hierarchical processing, self-attention mechanisms, and dynamic sparsity, enabling them to process large amounts of data efficiently. For example, Swin Transformers liu2021swin utilize a window-based self-attention mechanism, which reduces the computational complexity of the standard attention mechanism, making it feasible to scale models to much larger datasets and sequences. Proponents of this view argue that as these advanced models continue to evolve, there may be less need for aggressive input compression, as these models are inherently better equipped to handle large-scale data directly.
@@
Counterargument. Advanced model architectures offer strong performance but demand substantial computational resources, especially during training llama4 ; yang2025Qwen3 ; OpenAI:GPT-4 . Data-centric compression reduces computational load early by simplifying input data, enabling more efficient training and inference without sacrificing accuracy. Techniques like token pruning and augmentation preserve or improve performance by focusing on informative data. Combined with advanced architectures, data-centric methods enhance efficiency and maintain high performance gao2020pile ; wettig2024qurating , making them complementary rather than competitive.
7 Conclusion
In this position paper, we propose repositioning AI efficiency research by advocating a shift from model-centric to data-centric compression strategies, focusing on token compression to address long-context processing challenges. We first examine recent developments in long-context capabilities across various downstream scenarios, demonstrating how performance scaling has shifted from model size to context length, emphasizing the need for token compression to mitigate the impact of increasing context lengths. We then review approaches for improving model efficiency, with emphasis on the research roadmap of data-centric compression (i.e., token compression) and its potential benefits. After analyzing current challenges in token compression research, we propose promising future directions to inspire innovation in this emerging field. Our work aims to advance AI efficiency by providing a fresh perspective and catalyzing new research directions.
Appendix A Trends in LLM Scaling: Parameters vs. Context Length
In this section, we provide a comprehensive analysis of the temporal progression of mainstream LLMs, documenting the growth trends in both parameter counts and context lengths. This analysis provides empirical support for our central thesis regarding the shift in computational bottlenecks from model parameters to context processing. As shown in Tables 1, 2, 3, 4, 5, 6, 7, and 8, both in text and vision domains, model size growth has significantly slowed, while context length continues to increase. This trend indicates that the focus of research for efficient AI is shifting from model-centric compression to data-centric compression.
In this section, we conduct a detailed analysis comparing carefully designed token compression methods with random token dropping (the simplest baseline for token compression). This analysis aims to support the arguments presented in Section 4.1, demonstrating that existing token compression techniques have certain performance limitations. Our experiments span across multiple domains, including: complex reasoning in the language domain, image and video understanding in the vision domain, and text-to-image generation in the AI content generation domain. This comprehensive approach allows us to evaluate the effectiveness of token compression methods across diverse AI tasks and modalities. Our findings underscore the need for more robust and effective approaches in token compression for LLMs, MLLMs, VideoLLMs, and DiTs, highlighting the importance of developing universally applicable compression strategies.
LLMs: Complex Reasoning
We evaluated DeepSeek-R1-Distill-Llama-8B [55] on a suite of complex reasoning tasks, including MATH-500 [91], AIME24 [109], and GSM8K [31]. During the LLM’s decoding phase, we enforced a fixed token budget (e.g., 1024 tokens) for the KV cache and applied existing KV cache token dropping strategies such as H2O [172], SnapKV [88], KNorm [39], along with random dropping at regular intervals (e.g., every 512 tokens).
Figure 2 (a) reveals a counterintuitive finding: existing KV cache token dropping strategies—including H2O, SnapKV, and KNorm—consistently underperform compared to simple random token dropping across complex math reasoning tasks. Most strikingly, on AIME24, random dropping even surpasses the second-best method (SnapKV) by a significant margin of 10% accuracy. Our findings yield two critical insights for the field: (i) We strongly suggest that random dropping should be included as a fundamental baseline in KV cache dropping studies, as it is frequently overlooked in current research despite its competitive performance; (ii) We hypothesize that random dropping’s unexpected effectiveness may stem from its inherent property of preserving token distribution uniformity during auto-regressive decoding, thereby better maintaining semantic coherence and information integrity compared to deterministic dropping strategies. Our findings challenge the conventional wisdom that complex token dropping policies are inherently superior, while revealing fundamental gaps in current token importance modeling paradigms for KV cache management.
MLLMs: Image Understanding
We conducted experiments on multiple widely used image understanding benchmarks (e.g., GQA [65] and MMB [104]) using LLaVA-1.5-7B [98]. For all experiments, we uniformly retained 25% of the visual tokens, and benchmark evaluations were performed in accordance with the official evaluation scripts of LLaVA [^1] https://github.com/haotian-liu/LLaVA. In our experiments, we compared representative token compression methods, including FastV [19] and SparseVLM [170], along with two simple baselines: random token dropping and token-wise pooling.
As shown in Figure 2 (b), models employing random token dropping and token-wise pooling surprisingly outperform even some carefully designed methods. We hypothesize that the underlying reason lies in a key shared characteristic of random token dropping and token-wise pooling: spatial uniformity. This property effectively mitigates the issue of position bias (Sec. 4) inherent in attention-based token compression methods such as FastV. It also indirectly highlights the negative impact that position bias in attention scores can have on model performance. Therefore, we advocate for incorporating spatial uniformity as a key consideration in the design of token compression strategies.
VideoLLMs: Video Understanding
We conducted comprehensive comparative experiments on LLaVA-OneVision-7B [82] across multiple video large language models (VideoLLMs) benchmarks, including MVBench [86], LongVideoBench [148], MLVU [177], and VideoMME [44]. Our study compares various token compression methods, specifically FastV [19], SparseVLM [170], and PDrop [151], with the token retention ratio set to $R=15\%$. All evaluations were performed using the LMMs-Eval framework [166], ensuring consistency and reproducibility in our experimental setup [^2] https://github.com/LLaVA-VL/LLaVA-NeXT/blob/main/docs/LLaVA_OneVision.md.
Figure 2 (c) reveals an unexpected result: even when retaining only 15% of visual tokens, random token dropping outperforms carefully designed token compression methods (e.g., FastV [19], SparseVLM [170], and PDrop [151]). This finding has two important implications: (i) Future token compression studies should include random token dropping as a baseline for comparative experiments; (ii) The design of token compression methods for VideoLLMs should prioritize achieving a uniform spatial and temporal distribution of tokens, potentially ensuring a more comprehensive representation of video information. We hypothesize that the success of random dropping may be attributed to its inherent ability to maintain this uniformity across the video sequence.
DiTs: Image Generation
We conducted experiments on the widely-used DiT-based image generation model FLUX.1-dev [78], using the classic ToCa [181] method as an example. We set the cache cycle length to $N=4$ and the cache ratio to $R=90\%$, meaning that only 10% of the tokens were computed at each cache step. We compared several classic token selection strategies, such as those based on maximum Attention, Key Norm (i.e., Knorm), Value Norm (i.e., Vnorm), and random selection. Surprisingly, all of these characteristic-based selection strategies yielded lower Image Reward scores than random selection.
Figure 2 (d) shows that random selection performs well in token compression for image generation tasks. To investigate the reasons behind this phenomenon, we designed a similarity-based token selection strategy: we first randomly selected 1% of the tokens as base tokens, and then selected the remaining 9% based on their highest similarity to the base tokens. This selection process resulted in the chosen tokens being clustered into one or a few similar types, leading to the worst generation quality. This suggests that redundancy exists among similar tokens, whereas random selection benefits from diversity, allowing the selected subset of tokens to carry richer and more varied information.
Table 1: Qwen series model specifications. Includes release dates, parameter counts, maximum context lengths, and Hugging Face links.
@@ Appendix B Comparison of Token Compression Methods and Random Token Dropping
Table 2: DeepSeek series model specifications. Includes release dates, parameter counts, and maximum context lengths.
@@
Table 4: GLM series model specifications. Includes release dates, parameter counts, maximum context lengths, and Hugging Face links.
@@
Table 5: InternLM series model specifications. Includes release dates, parameter counts, maximum context lengths, and Hugging Face links.
@@
InternLM3-8B-Instruct January 15, 2025 8B 32768 tokens link
Table 6: LLaVA series model specifications. Includes release dates, backbone models, context lengths, and multimodal capabilities. si: single image; mi: multiple images; vid: video.
@@
Table 7: InternVL series model specifications. Includes release dates, backbone architectures, and multimodal capabilities.

### v3

@@ 4.2 Suboptimal Data Representation
Most existing data-centric compression methods fall into two categories: redundancy-based approaches that maximize information preservation between original ($\mathbf{X}$) and compressed data ($\mathbf{X^{\prime}}$) via $\max_{\mathcal{C}}I(\mathbf{X};\mathbf{X^{\prime}})$, and importance-based methods that ensure predictive sufficiency through $I(\mathbf{X^{\prime}};\mathbf{Y})\geq I(\mathbf{X};\mathbf{Y})-\epsilon$, where $\epsilon$ denotes bearable information loss. While effective for their respective objectives, we argue that these paradigms share a critical limitation: neither guarantees that the compressed data $\mathbf{X^{\prime}}$ forms an optimal representation for downstream modeling. The redundancy-based framework, despite preserving maximal mutual information with $\mathbf{X}$, often retains tokens with reconstructive but low discriminative value. The importance-based framework, on the other hand, prioritizes maintaining predictive performance with respect to the target variable $\mathbf{Y}$, but often at the cost of introducing task-specific biases. By focusing solely on information relevant to a predefined label, these methods may overlook the need to maintain stable structural and semantic patterns across the sequence that could enhance generalization. Consequently, both approaches risk producing representations misaligned with the ultimate goal of effective and generalizable downstream modeling.
@@ 4.3 Fair Comparison
Rethinking FLOPs and Compression Ratios as Efficiency Metrics. Many data-centric compression methods report speedup by estimating FLOPs reductions or directly using token compression ratios. But do FLOPs or compression ratios truly reflect real acceleration? Our analysis shows that, even with similar compression ratios or FLOPs, methods often vary significantly in runtime latency. Investigating further, we find: (i) Importance-based compression often uses attention scores Chen et al. (2024b); Zhang et al. (2025b), but this can limit compatibility with efficient attention operators (e.g., Flash Attention Dao (2024)), contributing to the discrepancy between theoretical FLOPs and actual latency. (ii) Some methods pursue high compression via progressive compression across layers Xing et al. (2025), adding computational overhead that offsets the gains from token reduction. Thus, we argue that runtime latency should be prioritized in evaluations, as FLOP or token count reductions do not always yield real-world speedup.
Data-centric Evaluation: The Benchmarking Gap. Current data-centric compression methods are mostly evaluated on general-purpose benchmarks that are not designed to capture compression-specific challenges. Consequently, benchmarks such as ScienceQA Lu et al. (2022) and VizWiz Bigham et al. (2010) sometimes show improved or stable performance under compression, contradicting expectations. This suggests these benchmarks may not adequately reflect the trade-offs inherent in compression. The issue undermines the reliability of current evaluations. Benchmarks insensitive to information loss can mask real differences between methods. Moreover, limited task diversity and the absence of compression-sensitive metrics hinder understanding of method behavior in practical settings. Without dedicated benchmarks, it is unclear whether reported gains reflect true progress or artifacts of misaligned evaluation.
@@ 5.1 Data-Model Centric Compression Co-Development
5.2 Dedicated Benchmarks for Data-centric Compression
@@
Given the current limitations in evaluating data-centric compression methods using general-purpose benchmarks, we envision the development of a dedicated benchmark specifically designed to evaluate them. Such a benchmark should comprehensively span diverse domains—including natural language processing, computer vision, and multi-modal tasks—and incorporate task-specific challenges particularly relevant to token compression, such as optical character recognition (OCR) parsing Ouyang et al. (2024); Zhang et al. (2024a) and automatic speech recognition (ASR) Conneau et al. (2023); Park et al. (2024). Furthermore, it is essential that this benchmark jointly considers both task performance and latency, as both are critical for real-world deployment. A well-rounded benchmark would enable a more rigorous, fair, and holistic evaluation of data-centric compression methods, ultimately driving progress in this area.
6 Conclusion
In this position paper, we propose repositioning AI efficiency research by advocating a shift from model-centric to data-centric compression to address long-context processing challenges. We first examine recent advances in long-context capabilities across downstream scenarios, showing that performance scaling has shifted from model size to context length, underscoring the need for data-centric compression to mitigate the overhead of growing context lengths. We then review model efficiency approaches, with emphasis on the research roadmap of data-centric compression and its potential benefits. After analyzing current challenges in this area, we outline promising future directions to inspire innovation. Our work aims to advance AI efficiency by offering a fresh perspective and catalyzing new research.
7 Limitations
In this work, we review data-centric compression methods and analyze their benefits and limitations. Due to space constraints, our analysis focuses on token overhead and compression techniques in several prominent domains (e.g., LLMs, MLLMs, and AIGC). We acknowledge that other application areas, such as computer vision, autonomous driving, embodied intelligence, and audio/speech processing, also face growing efficiency challenges and may benefit from data-centric compression. More comprehensive cross-domain review and analysis are left for our future work.
Appendix A Trends in LLM Scaling: Parameters vs. Context Length
In this section, we provide a comprehensive analysis of the temporal progression of mainstream LLMs, documenting the growth trends in both parameter counts and context lengths. This analysis provides empirical support for our central thesis regarding the shift in computational bottlenecks from model parameters to context processing. As shown in Tables below, both in text and vision domains, model size growth has significantly slowed, while context length continues to increase. This trend indicates that the focus of research for efficient AI is shifting from model-centric compression to data-centric compression.
This section presents a detailed comparison between designed token compression methods and random token dropping, the simplest baseline. The analysis supports arguments in Section 4.1, showing that current token compression techniques have performance limitations. Experiments span multiple domains: complex reasoning in language, image and video understanding in vision, and text-to-image generation in AI content creation. This broad evaluation assesses token compression effectiveness across diverse tasks and modalities. Results highlight the need for more robust approaches in LLMs, MLLMs, VideoLLMs, and DiTs, emphasizing the importance of universally applicable compression strategies.
LLMs: Complex Reasoning
We evaluated DeepSeek-R1-Distill-Llama-8B Guo et al. (2025a) on complex reasoning tasks: MATH-500 Lightman et al. (2023), AIME24 Maxwell-Jia (2024), and GSM8K Cobbe et al. (2021). During decoding, we enforced a fixed KV cache budget (i.e., 1024 tokens) and applied KV cache evition methods: H2O Zhang et al. (2023), SnapKV Li et al. (2024e), KNorm Devoto et al. (2024), and random eviction at regular intervals (i.e., every 512 tokens).
Figure 3 (a) shows a counterintuitive result: H2O, SnapKV, and KNorm consistently underperform random dropping on math reasoning tasks. On AIME24, random dropping surpasses SnapKV by 10% accuracy. These findings yield two insights: (i) Random dropping should be a standard baseline in KV cache studies, as it is often overlooked despite strong performance; (ii) Its effectiveness may stem from preserving token distribution uniformity during auto-regressive decoding, better maintaining semantic coherence than deterministic strategies. This challenges the assumption that complex policies are inherently superior and reveals gaps in current token importance modeling.
MLLMs: Image Understanding
We tested LLaVA-1.5-7B Liu et al. (2023) on image benchmarks including GQA Hudson and Manning (2019) and MMB Liu et al. (2025c). We retained 25% of visual tokens, following official LLaVA evaluation scripts [^1] https://github.com/haotian-liu/LLaVA. We compared FastV Chen et al. (2024b), SparseVLM Zhang et al. (2025b), random dropping, and pooling.
Figure 3 (b) shows that random dropping and pooling outperform some designed methods. We attribute this to their shared property of spatial uniformity, which mitigates position bias (Sec. 4) in attention-based methods like FastV. This also highlights the negative impact of position bias in attention scores. We thus advocate for spatial uniformity as a key design principle in token compression.
VideoLLMs: Video Understanding
We evaluated LLaVA-OneVision-7B Li et al. (2024a) on video benchmarks: MVBench Li et al. (2024c), LongVideoBench Wu et al. (2024), MLVU Zhou et al. (2024), and VideoMME Fu et al. (2024). We compared FastV Chen et al. (2024b), SparseVLM Zhang et al. (2025b), and PDrop Xing et al. (2025) with $15\%$ visual tokens, using the LMMs-Eval framework Zhang et al. (2024b) [^2] https://github.com/LLaVA-VL/LLaVA-NeXT/blob/main/docs/LLaVA_OneVision.md.
Figure 3 (c) shows that even with only 15% tokens, random dropping outperforms designed methods. This implies: (i) Random dropping must be included as a baseline; (ii) VideoLLM compression should prioritize uniform spatial and temporal token distribution for comprehensive video representation. We hypothesize that random dropping succeeds by inherently maintaining this uniformity.
DiTs: Image Generation
We tested the DiT-based model FLUX.1-dev Labs (2024) with the ToCa Zou et al. (2025) method, setting cache cycle length $N=4$ and ratio $R=90\%$ (10% tokens computed per step). We compared attention-based, Key Norm (Knorm), Value Norm (Vnorm), and random selection. Surprisingly, all characteristic-based strategies yielded lower Image Reward scores than random selection.
Figure 3 (d) confirms random selection performs well in image generation. To investigate, we designed a similarity-based strategy: select 1% tokens randomly as base, then choose 9% most similar to them. This led to clustered, homogeneous tokens and the worst generation quality. This suggests redundancy among similar tokens, while random selection benefits from diversity, enabling richer information representation.
Appendix C Alternative Positions
@@
C.1 Model-Centric Compression as a Superior Alternative
@@
Model-centric compression methods, such as pruning Liu et al. (2019); Huang et al. (2011); Han et al. (2016); Rao et al. (2021); Jang et al. (2017); Pan et al. (2021), quantization Rokh et al. (2023); Wang et al. (2019); Zhou et al. (2018); Yang et al. (2019); Lin et al. (2024a); Frantar et al. (2022), and knowledge distillation Hinton et al. (2015); Zhang et al. (2019, 2021); Park et al. (2019), have long been established as effective techniques for reducing model size and computational cost. Proponents argue that this paradigm is reliable for deployment in resource-constrained environments and maintains performance consistency. For example, pruning techniques such as DynamicViT Rao et al. (2021) dynamically remove uninformative tokens during inference, reducing the computational load by up to 30–40% with minimal impact on accuracy. Proponents of this view claim that this approach achieves substantial speedups without discarding any original data. In contrast, data-centric methods that prune input tokens risk removing critical contextual information, which may degrade performance.
@@
Counterargument. Although model-centric compression is effective, it faces scalability issues as models and datasets grow, requiring costly full retraining and processing of entire inputs. In contrast, data-centric compression reduces input complexity upfront, easing computational burdens. Some data-centric methods update only a small parameter subset Zhao et al. (2024b); Liu et al. (2024d); Lee and Hong (2024), while others enable training-free deployment Bolya et al. (2023a); Chen et al. (2024b); Li et al. (2024e). Combining both approaches can improve efficiency without sacrificing accuracy Azeemi et al. (2023); Kousar et al. (2025), making data-centric methods a complement to model-centric techniques.
C.2 Advanced Model Architectures as a more Promising Direction
@@
Another argument against data-centric compression is the continued advancement of model architectures that can inherently handle large datasets and long sequences more efficiently Gu et al. (2021); Goldstein et al. (2024); Peng et al. (2023). The development of transformer-based architectures, such as Vision Transformers Covert et al. (2022), Swin Transformers Liu et al. (2021), and large language models like GPT-3 Brown et al. (2020), has shown significant improvements in both accuracy and scalability. These architectures integrate advanced techniques, such as hierarchical processing, self-attention mechanisms, and dynamic sparsity, enabling them to process large amounts of data efficiently. For example, Swin Transformers Liu et al. (2021) utilize a window-based self-attention mechanism, which reduces the computational complexity of the standard attention mechanism, making it feasible to scale models to much larger datasets and sequences. Proponents of this view argue that as these advanced models continue to evolve, there may be less need for aggressive input compression, as these models are inherently better equipped to handle large-scale data directly.
@@
Counterargument. Advanced model architectures offer strong performance but demand substantial computational resources, especially during training Meta (2025); Yang et al. (2025a); OpenAI (2023). Data-centric compression reduces computational load early by simplifying input data, enabling more efficient training and inference without sacrificing accuracy. Techniques like token pruning and augmentation preserve or improve performance by focusing on informative data. Combined with advanced architectures, data-centric methods enhance efficiency and maintain high performance Gao et al. (2020); Wettig et al. (2024), making them complementary rather than competitive.
@@
Table 1: Qwen series model specifications. Details include release dates, parameter counts, maximum context lengths, and Hugging Face links.
@@
Table 2: DeepSeek series model specifications. Details include release dates, parameter counts, and maximum context lengths.
@@
Table 4: GLM series model specifications. Details include release dates, parameter counts, maximum context lengths, and Hugging Face links.
@@
InternLM3-8B-Instruct January 15, 2025 8B 32,768 tokens link
Table 5: InternLM series model specifications. Details include release dates, parameter counts, maximum context lengths, and Hugging Face links.
@@
Table 6: LLaVA series model specifications. Details include release dates, backbone models, context lengths, and multimodal capabilities. si: single image; mi: multiple images; vid: video.
@@
Table 7: InternVL series model specifications. Details include release dates, backbone architectures, and multimodal capabilities.